OpenStudy (anonymous):

Jimmy invests $2,000 at an interest rate of 10% for 4 years, while Jenny invests $2,000 at an interest rate of 5% for 8 years. Part 1: Determine the amount of return gained by Jimmy and Jenny. (3 points) Part 2: Summarize your results from Part 1, including how you arrived at your answer.

OpenStudy (anonymous):

Help? I think i did something wrong Jimmy 2000((1+(.1/12))^(12*4) = 2978.71 Jenny: 2000((1)+(.05/12))^(12*8)= 2981.17 using the equation a=p((1+(r/12))^(12t)

OpenStudy (phi):

the problem does not say if this is compound interest or simple interest (I= P*rate*time). You assumed compounded monthly. You should state so. In any case, the return gained is just the interest (not the principal) so subtract off the 2000.00

OpenStudy (anonymous):

Oh gosh thank you!

OpenStudy (anonymous):

Do you know if i did the math right though?